### v2

# A Garden of Eden theorem for Smale spaces

Abstract.
Given a dynamical system $(X,f)$ consisting of a compact metrizable space $X$ and a homeomorphism $f\colon X\to X$, an endomorphism of $(X,f)$ is a continuous map of $X$ into itself which commutes with $f$. One says that a dynamical system $(X,f)$ is surjunctive if every injective endomorphism of $(X,f)$ is surjective. An endomorphism of $(X,f)$ is called pre-injective if its restriction to each $f$-homoclinicity class of $X$ is injective. One says that a dynamical system has the Moore property if every surjective endomorphism of the system is pre-injective and that it has the Myhill property if every pre-injective endomorphism is surjective. One says that a dynamical system satisfies the Garden of Eden theorem if it has both the Moore and the Myhill properties. We prove that every irreducible Smale space satisfies the Garden of Eden theorem and that every non-wandering Smale space is surjunctive and has the Moore property.
Key words and phrases:
Smale space, Ruelle-Smale dynamical system, Myhill property, Moore property, Garden of Eden teorem, subshift of finite type, expansivity, shadowing, homoclinicity
2020 Mathematics Subject Classification:
37D20, 58F15, 54H20, 37B40, 37J45, 37B10, 37C50, 37B15, 37D20, 37C29
1. Introduction
Smale spaces were introduced by Ruelle at the end of the 1970s in his influential monograph on thermodynamic formalism [27]. They form a rich class of uniformly hyperbolic dynamical systems which includes in particular all Anosov diffeomorphisms [1], all non-wandering sets of Smale’s Axiom A diffeomorphisms [29], all expansive principal algebraic dynamical systems [28], [17], [20], all subshifts of finite type [18], as well as various types of solenoids and attractors [34], [33]. Here, by a dynamical system, we mean a pair $(X,f)$ consisting of a compact metrizable space $X$, called the phase space of the system, equipped with a homeomorphism $f\colon X\to X$. Roughly speaking, a dynamical system $(X,f)$ is a Smale space if $X$ admits a local product structure such that $f$ is uniformly contracting in one direction and uniformly expanding in the other one (see Section 2 for precise definitions). An endomorphism of a dynamical system $(X,f)$ is a continuous map $\tau\colon X\to X$ which commutes with $f$, i.e., such that $f\circ\tau=\tau\circ f$. One says that a dynamical system $(X,f)$ is surjunctive if every injective endomorphism of $(X,f)$ is surjective. As a trivial example, any dynamical system with finite phase space is surjunctive since every injective map of a finite set into itself is surjective. On the other hand, the dynamical system $(X,f)$, where $X$ is the real interval $[0,1]$ and $f\colon X\to X$ is the identity map, is not surjunctive since the map $\tau\colon X\to X$, defined by $\tau(x)\coloneqq x/2$ for all $x\in X$, is an injective endomorphism of $(X,f)$ which is not surjective. The word “surjunctive” was created by Gottschalk [15].
Given a dynamical system $(X,f)$ and a compatible metric on $X$, one says that two points $x,y\in X$ are homoclinic if the distance from $f^{n}(x)$ to $f^{n}(y)$ converges to zero as $|n|$ tends to infinity. This definition does not depend on the choice of the compatible metric by compactness of $X$. Homoclinicity is an equivalence relation on the phase space. One says that an endomorphism of a dynamical system is pre-injective if its restriction to every homoclinicity class of the phase space is injective. One says that a dynamical system $(X,f)$ has the Moore property if every surjective endomorphism of $(X,f)$ is pre-injective and that $(X,f)$ has the Myhill property if every pre-injective endomorphism of $(X,f)$ is surjective. Note that injectivity implies pre-injectivity. Therefore, every dynamical system which has the Myhill property is surjunctive. One says that a dynamical system has the Moore-Myhill property, or that it satisfies the Garden of Eden theorem, if it satisfies both the Moore and the Myhill properties. The Garden of Eden theorem was first established for full shifts by Moore [21] and Myhill [22]. It was subsequently extended to all irreducible subshifts of finite type by Fiorenzi in [14]. She also showed that irreducible sofic subshifts have the Myhill property but may fail to have the Moore property. In [8, Theorem 1.1], it is shown that Anosov diffeomorphisms on tori satisfy the Garden of Eden theorem.
In the last decades, several results on surjunctivity and Garden of Eden type theorems have been established in many other categories of dynamical systems (continuous actions of groups other than $\mathbb{Z}$, additional structure on the phase space, etc.). See the surveys [7], [9], the monographs [10], [11], and the references therein.
In the present paper, we shall prove the following.
Theorem 1.1.
Every irreducible Smale space satisfies the Garden of Eden theorem.
The result of Fiorenzi mentioned above, namely that irreducible subshifts of finite type satisfy the Garden of Eden theorem, is covered by Theorem 1.1 since all subshifts of finite type are Smale spaces. Theorem 1.1 also shows that every irreducible Anosov diffeomorphism satisfies the Garden of Eden theorem. As all Anosov diffeomorphisms on tori are known to be irreducible, Theorem 1.1 also covers Theorem 1.1 in [8]. However, it is an open question whether every Anosov diffeomorphism is irreducible. For an Anosov diffeomorphism $f\colon M\to M$, the following conditions are known to be equivalent (see e.g. [6, Theorem 5.10.3]): (a) $(M,f)$ is irreducible; (b) $(M,f)$ is topologically mixing; (c) $(M,f)$ is non-wandering; (d) the periodic points of $f$ are dense in $M$. All known examples of compact manifolds admitting Anosov diffeomorphisms are homeomorphic to infra-nilmanifolds, i.e., compact quotients of simply-connected nilpotent Lie groups by discrete groups of isometries. Manning [19] proved that all Anosov diffeomorphisms of infra-nilmanifolds are irreducible. Thus, Theorem 1.1 gives the following.
Corollary 1.2.
Every Anosov diffeomorphism of an infra-nilmanifold satisfies the Garden of Eden theorem.
There exist Smale spaces that are not surjunctive (see Remark 3.3) as well as Smale spaces that do not have the Myhill property nor the Moore property (see Remark 5.3). On the other hand, there exist non-wandering Smale spaces that do not have the Myhill property (Remark 5.2). However, we shall establish the following.
Theorem 1.3.
Every non-wandering Smale space is surjunctive and has the Moore property.
As every subshift of finite type is a Smale space, Theorem 1.3 gives us the following.
Corollary 1.4.
Let $A$ be a finite set. Then every non-wandering subshift of finite type $\Sigma\subset A^{\mathbb{Z}}$ is surjunctive and has the Moore property.
The paper is organized as follows. In Section 2, we introduce notation and collect background material. The proof of the first part of Theorem 1.3 is given in Section 3. The proof of Theorem 1.1 is given in Section 4. The proof of the second part of Theorem 1.3 is given in Section 5.
2. Preliminaries
In this section we fix notation and collect some background material.
2.1. General notation
We write $\mathbb{Z}$ for the set of integers and $\mathbb{N}$ for the set of non-negative integers.
Given a set $X$, we write $|X|$ for the cardinality of $X$ and we denote by $\operatorname{Id}_{X}$ the identity map on $X$. We use the symbol $\circ$ for the composition of maps (with the convention $(g\circ f)(x)\coloneqq g(f(x))$ for $f\colon X\to Y$, $g\colon Y\to Z$, and $x\in X$). Given a map $f\colon X\to X$, we set $f^{0}\coloneqq\operatorname{Id}_{X}$ and define inductively $f^{n}$ for $n\in\mathbb{N}$ by setting $f^{n+1}\coloneqq f\circ f^{n}$. In the case when $f$ is bijective, we write $f^{-1}$ for the inverse of $f$ and $f^{-n}\coloneqq(f^{-1})^{n}$ for every $n\in\mathbb{N}$.
If $Y$ is a subset of a topological space $X$, we write $\overline{Y}$ for the closure of $Y$ in $X$.
If $X$ is a metrizable space, we denote by $d_{X}$, or simply $d$ if there is no risk of confusion, any compatible metric on $X$.
2.2. Dynamical systems
(See [6], [27], [23])
Definition 2.1 (Dynamical system).
A dynamical system is a pair $(X,f)$, where $X$ is a compact metrizable space and $f\colon X\to X$ is a homeomorphism.
Let $(X,f)$ be a dynamical system. A subset $Y\subset X$ is said to be invariant if $f(Y)=Y$. If $Y\subset X$ is an invariant subset, we denote by $f|_{Y}$ the restriction of $f$ to $Y$, i.e., the homeomorphism $f|_{Y}\colon Y\to Y$ given by $f|_{Y}(y):=f(y)$ for all $y\in Y$. When $Y\subset X$ is a closed invariant subset, then the pair $(Y,f|_{Y})$ is also a dynamical system and one says that $(Y,f|_{Y})$ is a subsystem of $(X,f)$. We shall write it $(Y,f)$, or even simply $Y$, if there is no risk of confusion.
A point $x\in X$ is said to be periodic if there exists an integer $n\geq 1$ such that $f^{n}(x)=x$. One then says that the integer $n$ is a period of $x$. We denote by $\operatorname{Per}_{n}(X,f)$ the subset of $X$ consisting of all periodic points of $(X,f)$ admitting $n$ as a period. Note that $\operatorname{Per}_{n}(X,f)$ is a closed invariant subset of $X$. The set $\operatorname{Per}(X,f)\coloneqq\bigcup_{n\geq 1}\operatorname{Per}_{n}(X,f)$ of all periodic points of $f$ is also an invariant subset of $X$.
A point $x\in X$ is said to be non-wandering if for every neighborhood $U$ of $x$, there exists an integer $n\geq 1$ such that $U\cap f^{n}(U)\not=\varnothing$. The set $\operatorname{NW}(X,f)$ consisting of all non-wandering points of $f$ is a closed invariant subset of $X$. Observe that $\operatorname{Per}(X,f)\subset\operatorname{NW}(X,f)$. One says that the dynamical system $(X,f)$ is non-wandering if $\operatorname{NW}(X,f)=X$.
One says that $(X,f)$ is topologically mixing if for all non-empty subsets $U$ and $V$ of $X$, there exists an integer $N\geq 1$ such that $U\cap f^{n}(V)\not=\varnothing$ for all $n\geq N$. One says that $(X,f)$ is irreducible, or topologically $+$-transitive, if for all non-empty open subsets $U$ and $V$ of $X$, there exists $n\geq 1$ such that $U\cap f^{n}(V)\not=\varnothing$. The dynamical system $(X,f)$ is irreducible if and only if there exists a point $x\in X$ whose forward orbit $\{f^{n}(x):n\geq 1\}$ is dense in $X$. We have the following implications:
[…]
Each of these implications is strict. Indeed, consider a discrete space $X$ consisting of two distinct points and let $f\colon X\to X$ be the homeomorphism of $X$ which exchanges the two points of $X$. Then the dynamical system $(X,f)$ is irreducible but not topologically mixing. On the other hand, the dynamical system $(X,\operatorname{Id}_{X})$ is non-wandering but not irreducible.
Let $X$ be a compact metrizable space equipped with a homeomorphism $f\colon X\to X$. One says that the dynamical system $(X,f)$ has the weak specification property if for every $\varepsilon>0$ there exists an integer $N\geq 1$ such that for any sequence $I_{1},I_{2},\dots,I_{k}$ of finite intervals of $\mathbb{Z}$ satisfying $\min(I_{i+1})-\max(I_{i})\geq N$ for each $i\in\{1,2,\dots,k-1\}$ and any sequence $x_{1},x_{2},\dots,x_{k}$ of points of $X$, there exists a point $x\in X$ such that $d(f^{n}(x),f^{n}(x_{i}))\leq\varepsilon$ for all $i\in\{1,2,\dots,k\}$ and $n\in I_{i}$. By compactness of $X$, this definition does not depend on the choice of the compatible metric $d$.
Let $(X,f)$ and $(Y,g)$ be two dynamical systems. A continuous map $\varphi\colon X\to Y$ is called a morphism of dynamical systems if it satisfies $\varphi\circ f=g\circ\varphi$. Note that $\varphi(X)$ is then a closed invariant subset of $(Y,g)$. The class of all dynamical systems together with the morphisms between them form a category.
Proposition 2.2.
Let $\varphi\colon(X,f)\to(Y,g)$ be a morphism of dynamical systems. Then the following hold:
  1. (i)
    $\varphi(\operatorname{Per}_{n}(X,f))\subset\operatorname{Per}_{n}(Y,g)$ for every $n\geq 1$;
  2. (ii)
    $\varphi(\operatorname{Per}(X,f))\subset\operatorname{Per}(Y,g)$;
  3. (iii)
    $\varphi(\operatorname{NW}(X,f))\subset\operatorname{NW}(Y,g)$;
  4. (iv)
    if $(X,f)$ is non-wandering (resp. is irreducible, resp. is topologically mixing, resp. has the weak specification property) then $(\varphi(X),g)$ is non-wandering (resp. is irreducible, resp. is topologically mixing, resp. has the weak specification property).
Proof.
(i) If $x\in\operatorname{Per}_{n}(X,f)$, then $g^{n}(\varphi(x))=\varphi(f^{n}(x))=\varphi(x)$ so that $\varphi(x)\in\operatorname{Per}_{n}(Y,g)$.
(ii) Using (i), we get $\varphi(\operatorname{Per}(X,f))=\varphi(\bigcup_{n\geq 1}\operatorname{Per}_{n}(X,f))=\bigcup_{n\geq 1}\varphi(\operatorname{Per}_{n}(X,f))\subset\bigcup_{n\geq 1}\operatorname{Per}_{n}(Y,g)=\operatorname{Per}(Y,g)$.
(iii) Let $x\in\operatorname{NW}(X,f)$ and let $y\coloneqq\varphi(x)$. Let $V$ be a neighborhood of $y$ in $\varphi(X)$. Then $U\coloneqq\varphi^{-1}(V)$ is a neighborhood of $x$. Since $x\in\operatorname{NW}(X,f)$, there exists $n\geq 1$ such that $V\cap f^{n}(V)\not=\varnothing$. As $\varphi(V\cap f^{n}(V))\subset\varphi(V)\cap\varphi(f^{n}(V))=U\cap g^{n}(\varphi(V))\subset U\cap g^{n}(U)$, this implies $U\cap g^{n}(U)\not=\varnothing$. Therefore $y\in\operatorname{NW}(Y,g)$.
(iv) The fact that if $(X,f)$ is non-wandering then $(\varphi(X),g)$ is non-wandering immediately follows from (iii).
Observe that if $U$ and $V$ are any two non-empty open subsets of $\varphi(X)$, then $U^{\prime}\coloneqq\varphi^{-1}(U)$ and $V^{\prime}\coloneqq\varphi^{-1}(V)$ are two non-empty open subsets of $X$. As $U^{\prime}\cap f^{n}(V^{\prime})\not=\varnothing$ implies that $\varphi(U^{\prime}\cap f^{n}(V^{\prime}))\subset\varphi(U^{\prime})\cap\varphi(f^{n}(V^{\prime}))=\varphi(U^{\prime})\cap g^{n}(\varphi(V^{\prime}))\subset U\cap g^{n}(V)$, we deduce that if $(X,f)$ is irreducible (resp. topologically mixing) then $(\varphi(X),g)$ is irreducible (resp. topologically mixing).
Finally, suppose that $(X,f)$ has the weak specification property and let $\varepsilon>0$. Since $\varphi$ is uniformly continuous, there exists $\delta>0$ such that $d_{X}(x_{1},x_{2})\leq\delta$ implies $d_{Y}(\varphi(x_{1}),\varphi(x_{2}))\leq\varepsilon$ for all $x_{1},x_{2}\in X$. Suppose we are given a finite sequence $I_{1},I_{2},\dots,I_{k}$ of finite intervals of $\mathbb{Z}$ and a sequence $y_{1},y_{2},\dots,y_{k}$ of points of $\varphi(X)$. Choose $x_{1},x_{2},\dots,x_{k}\in X$ such that $\varphi(x_{i})=y_{i}$ for each $i\in\{1,2,\dots,k\}$. Since $(X,f)$ has the weak specification property, there exists an integer $N\geq 1$ such that if $\min(I_{i+1})-\max(I_{i})\geq N$ for each $i\in\{1,2,\dots,k-1\}$, then there exists $x\in X$ such that $d_{X}(f^{n}(x),f^{n}(x_{i}))\leq\delta$ for all $i\in\{1,2,\dots,k\}$ and $n\in I_{i}$. This implies that the point $y\coloneqq\varphi(x)\in\varphi(X)$ satisfies $d_{Y}(g^{n}(y),g^{n}(y_{i}))=d_{Y}(g^{n}(\varphi(x)),g^{n}(\varphi(x_{i})))=d_{Y}(\varphi(f^{n}(x)),\varphi(f^{n}(x_{i})))\leq\varepsilon$ for all $i\in\{1,2,\dots,k\}$ and $n\in I_{i}$. Thus, $(\varphi(X),g)$ has the weak specification property. This completes the proof of (iv). ∎
Two dynamical systems are said to be topologically conjugate if they are isomorphic objects in the category of dynamical systems. An isomorphism between two dynamical systems is also called a topological conjugacy. A surjective morphism $\pi\colon X\to Y$ is called a factor map. One says that $(X,f)$ is an extension of $(Y,g)$ and that $(Y,g)$ is a factor of $(X,f)$ if there exists a factor map $\pi\colon X\to Y$. It follows from Assertion (iv) in Proposition  2.2 that every factor of a dynamical system which is non-wandering (resp. irreducible, resp. topologically mixing, resp. with the weak specification property) is itself non-wandering (resp. irreducible, resp. topologically mixing, resp. with the weak specification property).
An endomorphism of the dynamical system $(X,f)$ is a continuous map $\tau\colon X\to X$ such that $\tau\circ f=f\circ\tau$. We write $\operatorname{End}(X,f)$ for the set of all endomorphisms of $(X,f)$. Observe that $f^{n}\in\operatorname{End}(X,f)$ for all $n\in\mathbb{Z}$ and that $\operatorname{End}(X,f)$ is a monoid for the composition of maps.
2.3. Stable equivalence, unstable equivalence, and homoclinicity
Let $(X,f)$ be a dynamical system and let $d$ be a compatible metric on $X$. Two points $x,y\in X$ are called stably equivalent (resp. unstably equivalent, resp. homoclinic), and one writes $x\sim_{s}y$ (resp. $x\sim_{u}y$, resp. $x\sim_{h}y$), if $d(f^{n}(x),f^{n}(y))\to 0$ as $n\to\infty$ (resp. $n\to-\infty$, resp. $|n|\to\infty$). Clearly, $\sim_{s}$ (resp. $\sim_{u}$, resp. $\sim_{h}$) is an equivalence relation on $X$. By compactness of $X$, each of these equivalence relations does not depend on the choice of the compatible metric $d$. Given $x\in X$, we denote by $W^{s}(X,f,x)$ (resp. $W^{u}(X,f,x)$, resp. $W^{h}(X,f,x)$), or simply $W^{s}(x)$ (resp. $W^{u}(x)$, resp. $W^{h}(x)$), the stable (resp. unstable, resp. homoclinic) equivalence class of $x$. Observe that $W^{h}(x)=W^{s}(x)\cap W^{u}(x)$ and that $f^{n}(W^{s}(x))=W^{s}(f^{n}(x))$ (resp. $f^{n}(W^{u}(x))=W^{u}(f^{n}(x))$, resp. $f^{n}(W^{h}(x))=W^{h}(f^{n}(x))$) for all $x\in X$ and $n\in\mathbb{Z}$.
Proposition 2.3.
Let $\varphi\colon(X,f)\to(Y,g)$ be a morphism of dynamical systems and let $x\in X$. Then the following hold:
  1. (i)
    $\varphi(W^{s}(X,f,x))\subset W^{s}(Y,g,\varphi(x))$;
  2. (ii)
    $\varphi(W^{u}(X,f,x))\subset W^{u}(Y,g,\varphi(x))$;
  3. (iii)
    $\varphi(W^{h}(X,f,x))\subset W^{h}(Y,g,\varphi(x))$.
Proof.
Let $d_{X}$ (resp. $d_{Y}$) be a compatible metric on $X$ (resp. $Y$). Take $x^{\prime}\in W^{s}(X,f,x)$. This means that $d_{X}(f^{n}(x^{\prime}),f^{n}(x))\to 0$ as $n\to\infty$. By continuity of $\varphi$, we have
$d_{Y}(\varphi(f^{n}(x^{\prime})),\varphi(f^{n}(x)))\to 0$
as $n\to\infty$. As $\varphi\circ f=g\circ\varphi$, we deduce that $d_{Y}(g^{n}(\varphi(x^{\prime})),g^{n}(\varphi(x)))\to 0$ as $n\to\infty$. This shows that $\varphi(x^{\prime})\in W^{s}(Y,g,\varphi(x))$ and hence that $\varphi(W^{s}(X,f,x))\subset W^{s}(Y,g,\varphi(x))$. The proofs of the two other inclusions are similar. It suffices to replace $n\to\infty$ by $n\to-\infty$ (resp. $|n|\to\infty$). ∎
Proposition 2.4.
Let $X$ be a compact metrizable space and let $f\colon X\to X$ be a homeomorphism. Let $k\geq 1$ be an integer and let $x,y\in X$. Then the following conditions are equivalent.
  1. (i)
    $x$ and $y$ are homoclinic points of $(X,f)$;
  2. (ii)
    $x$ and $y$ are homoclinic points of $(X,f^{k})$.
Proof.
Suppose first that $x$ and $y$ are homoclinic points of $(X,f)$. This means that $d(f^{n}(x),f^{n}(y))\to 0$ as $|n|\to\infty$. As this implies $d((f^{k})^{n}(x),(f^{k})^{n}(y))=d(f^{kn}(x),f^{kn}(y))\to 0$ as $|n|\to\infty$, we deduce that $x$ and $y$ are homoclinic for $(X,f^{k})$.
Conversely, suppose now that $x$ and $y$ are homoclinic points of $(X,f^{k})$. Let $\varepsilon>0$. By compactness of $X$, the maps $f,\dots,f^{k}$ are uniform homeomorphisms. Consequently, there exists $\delta>0$ such that $d(x_{1},x_{2})\leq\delta$ implies $d(f^{i}(x_{1}),f^{i}(x_{2}))\leq\varepsilon$ for all $x_{1},x_{2}\in X$ and $i\in\{1,\dots,k\}$. Since $x$ and $y$ are homoclinic for $(X,f^{k})$, there exists $N\geq 1$ such that $d(f^{kn}(x),f^{kn}(y))=d((f^{k})^{n}(x),(f^{k})^{n}(y))\leq\delta$ for all $n$ such that $|n|\geq N$. By our choice of $\delta$, this implies $d(f^{kn+i}(x),f^{kn+i}(y))=d(f^{i}(f^{kn}(x)),f^{i}(f^{kn}(y))\leq\varepsilon$ for all $n,i$ such that $|n|\geq N$ and $i\in\{1,\dots,k\}$. We deduce that $d(f^{n}(x),f^{n}(y))\leq\varepsilon$ for all $n$ such that $|n|\geq kN+1$. As $\varepsilon>0$ was arbitrary, this shows that the points $x$ and $y$ are homoclinic for $(X,f)$. ∎
One says that a morphism $\varphi\colon(X,f)\to(Y,g)$ of dynamical systems is pre-injective if the restriction of $\varphi$ to every homoclinicity class of $X$ is injective.
Proposition 2.5.
Let $\varphi\colon(X,f)\to(Y,g)$ and $\psi\colon(Y,g)\to(Z,h)$ be morphisms of dynamical systems. Suppose that $\psi\circ\varphi\colon(X,f)\to(Z,h)$ is pre-injective. Then $\varphi$ is pre-injective.
Proof.
Let $x_{1}$ and $x_{2}$ be homoclinic points in $X$ and suppose that $\varphi(x_{1})=\varphi(x_{2})$. We then have $(\psi\circ\varphi)(x_{1})=\psi(\varphi(x_{1}))=\psi(\varphi(x_{2}))=(\psi\circ\varphi)(x_{2})$. As $\psi\circ\varphi$ is pre-injective, this implies that $x_{1}=x_{2}$. This shows that $\varphi$ is pre-injective. ∎
Corollary 2.6.
Let $(X,f)$ be a dynamical system and let $\tau\in\operatorname{End}(X,f)$. Suppose that there exists an integer $k\geq 1$ such that $\tau^{k}$ is pre-injective. Then $\tau$ is pre-injective.
Proof.
This immediately follows from Proposition 2.5 after writing $\tau^{k}=\tau^{k-1}\circ\tau$. ∎
2.4. Expansivity
Let $X$ be a compact metrizable space and let $d$ be a compatible metric on $X$. A homeomorphism $f\colon X\to X$ is called expansive if there exists a constant $\eta>0$ such that if $x,y\in X$ and $x\not=y$ then there exists $n\in\mathbb{Z}$ such that $d(f^{n}(x),f^{n}(y))>\eta$. One then says that $\eta$ is an expansivity constant for $(X,f,d)$. By compactness, the definition of expansivity for a homeomorphism of $X$ does not depend on the choice of the metric $d$. Note that $\eta$ is an expansivity constant for $(X,f,d)$ if and only if it satisfies the following condition: for all $x,y\in X$, if one has $d(f^{n}(x),f^{n}(y))\leq\eta$ for all $n\in\mathbb{Z}$ then $x=y$.
Lemma 2.7.
local expansivity sauvegardeLet $X$ be a compact metrizable space and let $f\colon X\to X$ be an expansive homeomorphism. Then the set $\operatorname{Per}_{n}(X,f)$ is finite for every integer $n\geq 1$.
Proof.
Choose a compatible metric $d$ on $X$ and let $\eta>0$ be an expansivity constant for $(X,f,d)$. Let $n\geq 1$ be an integer. As the maps $f,f^{2},\dots,f^{n}$ are uniformly continuous, there exists $\varepsilon>0$ such that $d(f^{i}(x),f^{i}(y))\leq\eta$ for all $1\leq i\leq n$ and $x,y\in X$ with $d(x,y)\leq\varepsilon$. Suppose by contradiction that $\operatorname{Per}_{n}(X,f)$ is infinite. Then, by compactness of $X$, there are distinct points $x,y\in\operatorname{Per}_{n}(X,f)$ such that $d(x,y)\leq\varepsilon$. We then have $d(f^{i}(x),f^{i}(y))\leq\eta$ for all $i\in\mathbb{Z}$, a contradiction since $\eta$ is an expansivity constant for $(X,f,d)$. This shows that $\operatorname{Per}_{n}(X,f)$ is finite. ∎
2.5. Subshifts
(See [18], [11, Chapter 1]) Let $A$ be a finite discrete space. We equip the set $A^{\mathbb{Z}}$ of all bi-infinite sequences $u=(u_{n})_{n\in\mathbb{Z}}$, where $u_{n}\in A$ for all $n\in\mathbb{Z}$, with the topology of pointwise convergence. If we regard $A^{\mathbb{Z}}$ as the product of a family of copies of $A$ indexed by $\mathbb{Z}$, the topology on $A^{\mathbb{Z}}$ is the product topology. The space $A^{\mathbb{Z}}$ is compact, totally disconnected, and metrizable. A compatible metric $d$ on $A^{\mathbb{Z}}$ is defined by setting, for all $u,v\in A^{\mathbb{Z}}$,
$d(u,v)\coloneqq\begin{cases}0&\text{ if }u=v,\\ 2^{-\inf\{n\in\mathbb{N}:(u_{-n},u_{n})\not=(v_{-n},v_{n})\}}&\text{ if }u\not=v.\end{cases}$
The full shift over $A$ is the dynamical system $(A^{\mathbb{Z}},\sigma)$, where $\sigma\colon A^{\mathbb{Z}}\to A^{\mathbb{Z}}$ is the homeomorphism defined, for all $u\in A^{\mathbb{Z}}$, by $\sigma(u)\coloneqq v$ with $v_{n}\coloneqq u_{n-1}$ for all $n\in\mathbb{Z}$. Note that $(A^{\mathbb{Z}},\sigma)$ is expansive since $1/2$ is clearly an expansivity constant for $(A^{\mathbb{Z}},\sigma,d)$.
Two sequences $u,v\in A^{\mathbb{Z}}$ are $\sigma$-homoclinic if and only if there exists a finite subset $K\subset\mathbb{Z}$ such that $u_{n}=v_{n}$ for all $n\in\mathbb{Z}\setminus K$ [11, Exercise 5.2.(a)].
A closed $\sigma$-invariant subset $\Sigma\subset A^{\mathbb{Z}}$ is called a subshift.
Let $A^{*}$ denote the free monoid based on $A$. Every element of $A^{*}$ can be uniquely written as a word $w=a_{1}a_{2}\cdots a_{n}$, where $n\geq 0$ is the length of the word $w$, and $a_{i}\in A$ for all $1\leq i\leq n$. The monoid operation on $A^{*}$ is the concatenation of words. The language of a subshift $\Sigma\subset A^{\mathbb{Z}}$ is the subset $L(\Sigma)\subset A^{*}$ consisting of all words $w\in A^{*}$ such that there exist $n\geq 0$ and $u\in\Sigma$ such that $w=u_{1}u_{2}\cdots u_{n}$.
If $\Sigma\subset A^{*}$ is a subshift, one shows that the dynamical system $(\Sigma,\sigma)$ is irreducible if and only if, for all $u,v\in L(\Sigma)$, there exists $w\in A^{*}$ such that $uwv\in L(\Sigma)$ [11, Exercise 1.76].
A subshift $\Sigma\subset A^{\mathbb{Z}}$ is called a subshift of finite type if there exist an integer $m\geq 0$ and a subset $F\subset A^{*}$ consisting of words of length $m+1$ such that $\Sigma$ is the set of all sequences $u=(u_{n})_{n\in\mathbb{Z}}\in A^{\mathbb{Z}}$ satisfying $u_{n}u_{n+1}\cdots u_{n+m}\in F$ for all $n\in\mathbb{Z}$. One then says that $\Sigma$ is an $m$-step subshift of finite type and that $F$ is a defining set of admissible words for $\Sigma$. Observe that the full shift $A^{\mathbb{Z}}$ is a 00-step subshift of finite type admitting $A$ as a defining set of admissible words.
One says that a subshift $\Sigma\subset A^{\mathbb{Z}}$ is sofic if there exists a finite set $B$ such that $\Sigma$ is a factor of some subshift of finite type $\Sigma^{\prime}\subset B^{\mathbb{Z}}$.
A subshift $\Sigma\subset A^{\mathbb{Z}}$ is said to be strongly irreducible if there exists an integer $N\geq 1$ such that the following holds: if $I_{1}$ and $I_{2}$ are finite intervals of $\mathbb{Z}$ such that $\min(I_{2})-\max(I_{1})\geq N$ and $x_{1},x_{2}\in\Sigma$, then there exists $x\in\Sigma$ such that $x|_{I_{1}}=x_{1}|_{I_{1}}$ and $x|_{I_{2}}=x_{2}|_{I_{2}}$. It can be shown that a subshift $\Sigma\subset A^{\mathbb{Z}}$ has the weak specification property if and only if $\Sigma$ is strongly irreducible (cf. [17, Proposition A.1], [11, Exercise 1.69]). Every topologically mixing subshift of finite type is strongly irreducible (cf. [11, Exercise 1.80]). Consequently, we have the following.
Proposition 2.8.
Let $A$ be a finite set and let $\Sigma\subset A^{\mathbb{Z}}$ be a topologically mixing subshift of finite type. Then $\Sigma$ has the weak specification property.
2.6. The pseudo-orbit tracing property
Let $(X,f)$ be a dynamical system and let $d$ be a compatible metric on $X$.
Given $\varepsilon>0$, one says that a sequence $(u_{n})_{n\in\mathbb{Z}}$ of points of $X$ is $\varepsilon$-traced by the orbit of a point $x\in X$ if one has $d(u_{n},f^{n}(x))\leq\varepsilon$ for all $n\in\mathbb{Z}$.
Given $\delta>0$, one says that a sequence $(u_{n})_{n\in\mathbb{Z}}$ of points of $X$ is a $\delta$-pseudo-orbit if one has $d(u_{n+1},f(u_{n}))\leq\delta$ for all $n\in\mathbb{Z}$.
One says that the dynamical system $(X,f)$ has the pseudo-orbit tracing property if for every $\varepsilon>0$, there exists $\delta>0$ such that every $\delta$-pseudo-orbit in $X$ is $\varepsilon$-traced by the orbit of some point in $X$. By compactness of $X$, the fact that $f$ has the pseudo-orbit tracing property or not does not depend on the choice of the compatible metric $d$.
2.7. Smale spaces
We first recall Ruelle’s original definition of a Smale space (cf. [27, Section 7.1], [25, Definition 2.1.6] [26, Section 2.2]). Let $X$ be a compact metrizable space equipped with a homeomorphism $f\colon X\to X$. One says that $(X,f)$ is a Smale space if there exist a compatible metric $d$ on $X$, constants $\varepsilon>0$ and $10<\lambda<$ 1, and a continuous map
$[\cdot,\cdot]\colon\Delta_{\varepsilon}\coloneqq\{(x,y)\in X\times X:d(x,y)\leq\varepsilon\}\to X$
satisfying the following conditions for all $x,y,z\in X$:
  1. (Sm1)
    $[x,x]=x$ for all $x\in X$;
  2. (Sm2)
    $[[x,y],z]=[x,z]$ whenever both sides are defined;
  3. (Sm3)
    $[x,[y,z]]=[x,z]$ whenever both sides are defined;
  4. (Sm4)
    $[f(x),f(y)]=f([x,y])$ whenever both sides are defined;
  5. (Sm5)
    $d(f(y),f(z))\leq\lambda d(y,z)$ whenever $(y,x),(z,x)\in\Delta_{\varepsilon}$ and $[y,x]=x=[z,x]$;
  6. (Sm6)
    $d(f^{-1}(y),f^{-1}(z))\leq\lambda d(y,z)$ whenever $(x,y),(x,z)\in\Delta_{\varepsilon}$ and $[x,y]=x=[x,z]$.
It is known that every Smale space is expansive [27, Section 7.3], [25, Proposition 2.1.9] and has the pseudo-orbit tracing property [27, Section 7.3]. Conversely, if $(X,f)$ is a dynamical system such that $f$ is expansive and has the pseudo-orbit tracing property, then $(X,f)$ is a Smale space [24, Theorem]. Thus, an equivalent but more concise definition of Smale spaces is the following.
Definition 2.9 (Smale space).
A Smale space is a dynamical system $(X,f)$ consisting of a compact metrizable space $X$ equipped with an expansive homeomorphism $f\colon X\to X$ which has the pseudo-orbit tracing property.
The following result is known as Anosov’s closing lemma (cf. [5, 3.8], [27, Section 7.3]). It is a key ingredient in our proof of the first part of Theorem 1.3.
Theorem 2.10 (Anosov closing lemma for Smale spaces).
Let $(X,f)$ be a Smale space. Then the set $\operatorname{Per}(X,f)$ is dense in $\operatorname{NW}(X,f)$. In other words, we have $\overline{\operatorname{Per}(X,f)}=\operatorname{NW}(X,f)$.
Walters [30, Theorem 1] proved that a subshift $\Sigma\subset A^{\mathbb{Z}}$ has the pseudo-orbit tracing property if and only if $\Sigma$ is of finite type. As full shifts are expansive and every subsystem of an expansive dynamical system is itself expansive, it follows that a subshift $\Sigma\subset A^{\mathbb{Z}}$ is a Smale space if and only if $\Sigma$ is of finite type. Moreover, if $(X,f)$ is a Smale space with $X$ totally disconnected, then there exist a finite set $A$ and a $1$-step subshift of finite type $\Sigma\subset A^{\mathbb{Z}}$ such that $(X,f)$ is topologically conjugate to $(\Sigma,\sigma)$.
2.8. Coding of topologically mixing Smale spaces
The theory of Markov partitions used by Rufus Bowen in the setting of basic sets of Axiom A diffeomorphisms directly extends to Smale spaces an yields in particular the following result (see [27, Theorem 7.6], [2], [4], [5, Theorem 3.18 and Proposition 3.19]). It is a key ingredient in our proof of Theorem 1.1.
Theorem 2.11.
Let $(X,f)$ be a topologically mixing Smale space. Then there exist a finite set $A$ and a topologically mixing subshift of finite type $\Sigma\subset A^{\mathbb{Z}}$ such that $(X,f)$ is a factor of $(\Sigma,\sigma)$.
2.9. The Smale-Bowen-Ruelle decomposition theorem
The following result is known as the Smale-Bowen-Ruelle spectral decomposition theorem for non-wandering Smale spaces (see [5, Section 3.B], [27, Section 7.4], [26, Section 4.5]).
Theorem 2.12 (The Smale-Bowen-Ruelle decomposition theorem).
Let $(X,f)$ be a non-wandering Smale space. Then the following hold:
  1. (SBR1)
    the set $\mathcal{W}\coloneqq\{\overline{W^{u}(p)}:p\in\operatorname{Per}(X,f)\}$ is finite;
  2. (SBR2)
    the elements of $\mathcal{W}$ are non-empty clopen subsets of $X$ and form a finite partition of $X$;
  3. (SBR3)
    one has $f(W)\in\mathcal{W}$ for every $W\in\mathcal{W}$;
  4. (SBR4)
    the map $\mathcal{W}\to\mathcal{W}$, $W\mapsto f(W)$, is a permutation of $\mathcal{W}$;
  5. (SBR5)
    if $W\in\mathcal{W}$ and $k_{W}$ is the least positive integer such that $f^{k_{W}}(W)=W$, then the dynamical system $(W,f^{k_{W}})$ is a topologically mixing Smale space;
  6. (SBR6)
    for every $W\in\mathcal{W}$, the set $\Omega_{W}\coloneqq\bigcup_{0\leq i\leq k_{W}-1}f^{i}(W)$ is a non-empty $f$-invariant clopen subset of $X$ and the dynamical system $(\Omega_{W},f)$ is an irreducible Smale space;
  7. (SBR7)
    if $Y$ is a non-empty $f$-invariant clopen subset of $X$ such that $(Y,f)$ is irreducible, then there exists $W\in\mathcal{W}$ such that $Y=\Omega_{W}$.
Lemma 2.13.
Let $(X,f)$ be a non-wandering Smale space and let $\mathcal{W}$ be as in Theorem 2.12. Suppose that $\tau\in\operatorname{End}(X,f)$ and let $W\in\mathcal{W}$. Then there exists $W^{\prime}\in\mathcal{W}$ such that $\tau(W)\subset W^{\prime}$.
Proof.
Let $p\in\operatorname{Per}(X,f)$ such that $W=\overline{W^{u}(p)}$. We have $\tau(W^{u}(p))\subset W^{u}(\tau(p))$ by Proposition 2.3.(ii). Using the continuity of $\tau$, we deduce that $\tau(W)=\tau(\overline{W^{u}(p)})\subset\overline{\tau(W^{u}(p))}\subset\overline{W^{u}(\tau(p))}$. As $\tau(p)\in\operatorname{Per}(X,f)$ by Proposition 2.2.(ii), we deduce that we can take $W^{\prime}\coloneqq\overline{W^{u}(\tau(p))}\in\mathcal{W}$. ∎
2.10. The Li-Doucha Garden of Eden theorem
The following result was proved by Li [17] and Doucha [13], using properties of topological entropy, in the more general setting of expansive continuous actions with the weak specification property and the pseudo-orbit tracing property of countable amenable groups. Actually, the Myhill property for such systems was first obtained by Li [17, Theorem 1.1] and, a few years after, Doucha [13, Theorem A] established the Moore property. Note that the hypothesis that the system satisfies the pseudo-orbit tracing property is not needed for the Myhill part.
Theorem 2.14.
(Li [17], Doucha [13]) Let $(X,f)$ be a dynamical system consisting of a compact metrizable space $X$ and a homeomorphism $f\colon X\to X$. Suppose that $(X,f)$ is expansive and has both the weak specification property and the pseudo-orbit tracing property. Then $(X,f)$ satisfies the Garden of Eden theorem.
3. Surjunctivity of non-wandering Smale spaces
In this section, we prove the first part of Theorem 1.3.
The following auxiliary result is a particular case of Proposition 5.1 in [8].
Lemma 3.1.
Let $X$ be a compact metrizable space and let $f\colon X\to X$ be a homeomorphism. Suppose that the dynamical system $(X,f)$ is expansive and that the set $\operatorname{Per}(X,f)$ is dense in $X$. Then $(X,f)$ is surjunctive.
Proof.
Let $\tau\colon X\to X$ be an injective endomorphism of $(X,f)$. Let $n\geq 1$ be an integer. By Lemma 2.7, the set $\operatorname{Per}_{n}(X,f)$ is finite. As $\tau(\operatorname{Per}_{n}(X,f))\subset\operatorname{Per}_{n}(X,f)$ by Proposition 2.2.(a), injectivity of $\tau$ implies that $\tau(\operatorname{Per}_{n}(X,f))=\operatorname{Per}_{n}(X,f)$. Since $\operatorname{Per}(X,f)=\bigcup_{n\geq 1}\operatorname{Per}_{n}(X,f)$, it follows that $\tau(\operatorname{Per}(X,f))=\operatorname{Per}(X,f)$. The set $\tau(X)$ is closed in $X$ by continuity of $\tau$ and compactness of $X$. Since $\operatorname{Per}(X,f)$ is dense in $X$ by our hypotheses, we conclude that $\tau(X)=X$. This shows that $\tau$ is surjective and hence that $(X,f)$ is surjunctive. ∎
Theorem 3.2.
Every non-wandering Smale space is surjunctive.
Proof.
As every Smale space is expansive, this immediately follows from Theorem 2.10 and Lemma 3.1. ∎
Remark 3.3.
There exist non-surjunctive Smale spaces. In fact, there are subshifts of finite type that are not surjunctive. The following example was given by Weiss [32, p. 358] (cf. [11, Exercise 3.38]). Consider the $1$-step subshift of finite type $\Sigma\subset\{0,1,2\}^{\mathbb{Z}}$ admitting the set $\{00,11,22,01,12\}$ as a defining set of admissible words. Observe that the word $12$ can appear at most once in a configuration of $\Sigma$. It is easy to check that the map $\tau\colon\Sigma\to\Sigma$, $x\mapsto\tau(x)$, which replaces the word $12$ (if it appears in $x$) by $11$, is an injective endomorphism of $\Sigma$ which is not surjective (note that $012\in L(\Sigma)\setminus L(\tau(\Sigma))$). Therefore $(\Sigma,\sigma)$ is not surjunctive.
Remark 3.4.
A surjective endomorphism of a Smale space $(X,f)$ may fail to be injective even if $(X,f)$ is topologically mixing. For example, the endomorphism $\tau$ of the full shift ($(\{0,1\}^{\mathbb{Z}},\sigma)$, defined by $(\tau(u))_{n}\coloneqq u_{n}+u_{n+1}\mod 2$, is easily shown to be surjective but not injective (cf. [10, Example 3.3.8]). Another example is provided by Arnold’s cat, i.e., the Anosov diffeomorphism $f$ of the $2$-torus $X\coloneqq\mathbb{R}^{2}/\mathbb{Z}^{2}$ defined by $f(x_{1},x_{2})\coloneqq(2x_{1}+x_{2},x_{1}+x_{2})$, and the endomorphism $\tau$ of $(X,f)$ given by $\tau(x_{1},x_{2})\coloneqq(2x_{1},2x_{2})$ (cf. [8, Section 4]).
4. The Garden of Eden theorem for irreducible Smale spaces
The goal of this section is to establish Theorem 1.1. We shall use the following auxiliary results.
Lemma 4.1.
Let $(X,f)$ be a non-wandering Smale space and let $\mathcal{W}$ be as in Theorem 2.12. Suppose that $x,y\in X$ are stably equivalent (resp. unstably equivalent, resp. homoclinic) and let $W$ be the unique element of $\mathcal{W}$ such that $x\in W$. Then one has $y\in W$.
Proof.
Suppose that $x,y\in X$ are stably equivalent, i.e., $d(f^{n}(x),f^{n}(y))\to 0$ as $n\to\infty$. Let $k=k_{W}\geq 1$ be the integer as in Theorem 2.12.(SBR5), so that $f^{k}(W)=W$. Since $W$ is a compact open subset of $X$, there exists $\varepsilon>0$ such that $B_{X}(w,\varepsilon)\coloneqq\{w^{\prime}\in X:d(w,w^{\prime})<\varepsilon\}\subset W$ for every $w\in W$. As $x$ and $y$ are stably equivalent, there exists $n\geq 1$ such that $d(f^{kn}(x),f^{kn}(y))<\varepsilon$. Since $W$ is $f^{kn}$-invariant, this implies $f^{kn}(y)\in B_{X}(f^{kn}(x),\varepsilon)\subset W$ and hence $y\in W$.The proof in the case when $x$ and $y$ are unstably equivalent is similar. As homoclinic points are stably equivalent, the case when $x$ and $y$ are homoclinic follows as well. ∎
Lemma 4.2.
Every topologically mixing Smale space satisfies the Garden of Eden theorem.
Proof.
Let $(X,f)$ be a topologically mixing Smale space. By Theorem 2.11, there exist a finite set $A$ and a topologically mixing subshift of finite type $\Sigma\subset A^{\mathbb{Z}}$ such that $(X,f)$ is a factor of $(\Sigma,\sigma)$. As $(\Sigma,\sigma)$ has the weak specification property by Proposition 2.8, it follows from Proposition 2.2.(iv) that $(X,f)$ has also the weak specification property (cf. [27, Section 7.14]). Since every Smale space is expansive and has the pseudo-orbit tracing property, we deduce from Theorem 2.14 that $(X,f)$ satisfies the Garden of Eden theorem. ∎
Proof of Theorem 1.1.
Let $(X,f)$ be an irreducible Smale space. By the Smale-Bowen-Ruelle decomposition theorem (Theorem 2.12), there exist $k\in\mathbb{N}$ and a finite partition $X=\bigsqcup_{i\in\mathbb{Z}/k\mathbb{Z}}W_{i}$ into non-empty clopen subsets $W_{i}\subset X$, $i\in\mathbb{Z}/k\mathbb{Z}$, such that $f(W_{i})=W_{i+1}$ and $(W_{i},f^{k})$ is topologically mixing for every $i\in\mathbb{Z}/k\mathbb{Z}$.
Let $\tau\in\operatorname{End}(X,f)$. By Lemma 2.13, there is a unique map $\alpha\colon\mathbb{Z}/k\mathbb{Z}\to\mathbb{Z}/k\mathbb{Z}$ such that $\tau(W_{i})\subset W_{\alpha(i)}$ for all $i\in Z/k\mathbb{Z}$. We have that $\tau(W_{i+1})=\tau(f(W_{i}))=f(\tau(W_{i}))\subset f(W_{\alpha(i)})=W_{\alpha(i)+1}$ so that $\alpha(i+1)=\alpha(i)+1$ for all $i\in\mathbb{Z}/k\mathbb{Z}$. By induction, we get $\alpha(i)=i+\alpha(0)$ for all $i\in\mathbb{Z}/k\mathbb{Z}$. This implies in particular that $\alpha$ is a permutation of $\mathbb{Z}/k\mathbb{Z}$.
Let us choose, for each $i\in\mathbb{Z}/k\mathbb{Z}$, an integer $n_{i}\in\mathbb{Z}$ such that $i-\alpha(i)=n_{i}+k\mathbb{Z}$. Note that the map $\tau_{i}\colon W_{i}\to W_{i}$, defined by $\tau_{i}(x)\coloneqq f^{n_{i}}(\tau(x))$ for all $x\in W_{i}$, is an endomorphism of $(W_{i},f^{k})$.
Suppose first that $\tau$ is surjective. Then we must have $\tau(W_{i})=W_{\alpha(i)}$ for all $i\in\mathbb{Z}/k\mathbb{Z}$. Suppose that $x,y\in X$ are homoclinic for $(X,f)$ with $\tau(x)=\tau(y)$. Let $i$ be the unique element in $\mathbb{Z}/k\mathbb{Z}$ such that $x\in W_{i}$. Since $x$ and $y$ are homoclinic, it follows from Lemma 4.1 that $y\in W_{i}$. Consequently, we have $\tau(x),\tau(y)\in W_{\alpha(i)}$. Since the dynamical system $(W_{i},f^{k})$ is a topologically mixing Smale space, it has the Moore property by Lemma 4.2. As $\tau_{i}$ is a surjective endomorphism of $(W_{i},f^{k})$, we deduce that $\tau_{i}$ is pre-injective. Since $x$ and $y$ are homoclinic points of $(W_{i},f^{k})$ (by (i) $\implies$ (ii) in Proposition 2.4) with the same image under $\tau_{i}$, it follows that $x=y$. This shows that $\tau$ is pre-injective and hence that $(X,f)$ has the Moore property.
Suppose now that $\tau$ is a pre-injective endomorphism of $(X,f)$. Let $i\in\mathbb{Z}/k\mathbb{Z}$. If $x,y\in W_{i}$ are homoclinic for $(W_{i},f^{k})$ then they are homoclinic for $(X,f^{k})$ and hence for $(X,f)$ (by using (ii) $\implies$ (i) in Proposition 2.4). If, in addition, $\tau_{i}(x)=\tau_{i}(y)$ then $\tau(x)=\tau(y)$ so that $x=y$ by pre-injectivity of $\tau\in\operatorname{End}(X,f)$. Consequently, the map $\tau_{i}\colon W_{i}\to W_{i}$ is a pre-injective endomorphism of $(W_{i},f^{k})$. As $(W_{i},f^{k})$ has the Myhill property by Lemma 4.2, we deduce that $\tau_{i}(W_{i})=W_{i}$. This implies that $\tau(W_{i})=W_{\alpha(i)}$. As $\alpha$ is a permutation of $\mathbb{Z}/k\mathbb{Z}$, we conclude that $\tau(X)=X$. This shows that $\tau$ is surjective and hence that $(X,f)$ has the Moore property. ∎
Remark 4.3.
There is also a purely measure-theoretic characterization of surjectivity for endomorphisms of irreducible Smale spaces. Indeed, consider an irreducible Smale space $(X,f)$. By the work of Bowen (see [2], [3], [16]), the dynamical system $(X,f)$ is intrinsically ergodic in the sense of [31]. This means that the topological entropy $h(X,f)$ of $(X,f)$ is finite and there is a unique $f$-invariant Borel probability measure $\mu$ on $X$ whose measure-theoretic entropy $h_{\mu}(X,f)$ is equal to $h(X,f)$. Moreover, the support of $\mu$ is the entire space $X$. It then follows from Corollary 2.3 in [12] that an endomorphism $\tau$ of $(X,f)$ is surjective if and only if $\tau$ preserves $\mu$, i.e., if and only if $\mu(\tau^{-1}(B))=\mu(B)$ for all Borel subsets $B\subset X$.
5. The Moore property for non-wandering Smale spaces
In this section, we prove the second part of Theorem 1.3.
Theorem 5.1.
Every non-wandering Smale space has the Moore property.
Proof.
Let $(X,f)$ be a non-wandering Smale space and let $\tau\in\operatorname{End}(X,f)$ be a surjective endomorphism of $(X,f)$. Let us show that $\tau$ is pre-injective. We keep the notation as in Theorem 2.12. It follows from Lemma 2.13 that for every $W\in\mathcal{W}$ there exists a unique $\alpha(W)\in\mathcal{W}$ such that $\tau(W)\subset\alpha(W)$. As $\tau$ is surjective, the map $\alpha\colon\mathcal{W}\to\mathcal{W}$ is a permutation of $\mathcal{W}$ and we have $\tau(W)=\alpha(W)$ for every $W\in\mathcal{W}$. If $k$ denotes the order of the permutation $\alpha$ (i.e., the least integer $k\geq 1$ such that $\alpha^{k}=\operatorname{Id}_{\mathcal{W}}$), we have $\tau^{k}(W)=\alpha^{k}(W)=W$ for every $W\in\mathcal{W}$. Consequently, $\tau^{k}(\Omega_{W})=\Omega_{W}$ for every $W\in\mathcal{W}$. As the dynamical system $(\Omega_{W},f)$ is an irreducible Smale space, it has the Moore property by Theorem 1.1. As $\tau^{k}|_{\Omega_{W}}\in\operatorname{End}(\Omega_{W},f)$ is surjective, this implies that $\tau^{k}|_{\Omega_{W}}$ is pre-injective. Two homoclinic points in $X$ necessarily belong to the same $W\in\mathcal{W}$ (cf. Lemma 4.1) and therefore to the same $\Omega_{W}$. We deduce that $\tau^{k}\in\operatorname{End}(X,f)$ is pre-injective. It then follows from Corollary 2.6 that $\tau$ itself is pre-injective. This shows that $(X,f)$ has the Moore property. ∎
Remark 5.2.
There exist non-wandering Smale spaces which do not have the Myhill property. Consider for example the dynamical system $(X,f)$, where $X=\{x_{0},x_{1}\}$ is a discrete space with two distinct points and $f\coloneqq\operatorname{Id}_{X}$. It is clear that $(X,f)$ is a non-wandering Smale space with two homoclinicity classes, namely, $\{x_{0}\}$ and $\{x_{1}\}$. The map $\tau\colon X\to X$ defined by $\tau(x_{0})=\tau(x_{1})\coloneqq x_{0}$ is a pre-injective endomorphism of $(X,f)$ which is not surjective. Therefore, $(X,f)$ does not have the Myhill property. Observe that the dynamical system $(X,f)$ is topologically conjugate to the $1$-step subshift of finite type $\Sigma\subset\{0,1\}^{\mathbb{Z}}$ consisting of the two constant configurations.
Remark 5.3.
There exist Smale spaces which have neither the Moore property nor the Myhill property. Indeed, consider the Weiss subshift $\Sigma\subset\{0,1,2\}^{\mathbb{Z}}$ as in Remark 3.3. We have already observed that $\Sigma$ is not surjunctive, so that, a fortiori, it does not have the Myhill property. On the other hand, it is easy to see that the map $\tau\colon\Sigma\to\Sigma$, $x\mapsto\tau(x)$, which replaces the word $112$ (if it appears in $x$) by $122$, is a surjective endomorphism of $\Sigma$ which is not pre-injective (the sequences $u,v\in\Sigma$ defined by $u_{n}=v_{n}\coloneqq 0$ for all $n\leq 0$, $u_{1}=v_{1}=v_{2}\coloneqq 1$, $u_{2}=2$, and $u_{n}=v_{n}\coloneqq 2$ for all $n\geq 2$, are homoclinic and satisfy $\tau(u)=\tau(v)=u$). Therefore $\Sigma$ does not have the Moore property either.